# Amortized Active Generation of Pareto Sets

### v2

@@ 1 Introduction
Traditional Bayesian optimization (BO) methods for black-box multi-objective optimization (MOO) problems, often referred to as multi-objective Bayesian optimization (MOBO), rely on acquisition functions such as expected hypervolume improvement (EHVI) [38, 7, 1] or their quasi-MonteCarlo extensions [8, 1], which involve complex numerical integration and/or scale poorly with the number of objectives. Alternatively they may rely on random scalarizations [24, 28, 10] offering simplicity and scalability but rely on sufficient sampling density to capture complex Pareto front geometries.
In this work we propose a fundamentally different approach, in the spirit of multi-objective generation (MOG) [40, 39], that directly estimates a generative model of the Pareto set in an online sequential, black-box optimization setting. To this end, we build upon the recently proposed variational search distributions (VSD) framework for single-objective optimization [35]. VSD formulates optimization as active learning of a generative model of high-performing designs (active generation). Thus, VSD alternates between fitting a class probability estimator (CPE) to discriminate favorable designs and updating a conditional generative model to propose new candidates directly. Hence, instead of using a fitness-threshold probability of improvement (PI)-CPE, we propose a Pareto-set CPE, which enables the generative model to focus exclusively on non-dominated designs. This approach bypasses explicit hypervolume computations and scalarizations and leads to a scalable online algorithm for Pareto set approximation that is different from other recent MOG methods [40, 39]. Furthermore, we show that such a non-dominance CPE is implicitly estimating probability of hypervolume improvement (PHVI). Additionally, practical decision-making often requires incorporating user or stakeholder preferences over the trade-offs among objectives. We introduce a novel mechanism for embedding subjective preference conditioning into the generative model, allowing practitioners to sample candidates that align with specified trade-off directions. By using amortized variational inference (VI) with preference direction vectors, our method supports a-posteriori preference specification without retraining, offering flexibility in downstream design exploration.
@@ 2.1 Active generation
Active generation as implemented by [35] reframes online black-box optimization as sequential learning of a conditional generative model, guided by a CPE. At each round, $t\in\{1,\dots,T\}$ we: (1) fit a CPE (using some proper loss, $\mathcal{L}_{\text{CPE}}$),
@@
$\theta^{*}_{t}\leftarrow\operatorname*{arg\,\!min}_{\theta}\mathcal{L_{\textrm{CPE}}}\!\left({\theta,\mathcal{D}^{z}_{N}}\right)\qquad\phi^{*}_{t}\leftarrow\operatorname*{arg\,\!max}_{\phi}\mathcal{L}_{\text{ELBO}}(\phi,\theta^{*}_{t}),$ (3)
then samples from $q_{\phi^{*}_{t}}({\mathbf{x}})$ are used to propose new candidates for evaluation. New labels are acquired for these candidates, the dataset is augmented, and the process is repeated until convergence. This solution to active generation is referred to as VSD [35]. Under certain assumptions on the form of the models, this procedure has proven convergence rates to the ideal $p({\mathbf{x}}|{z=1})$.
@@ 2.2 Optimizing over multiple objectives
$\mathbf{x}^{\prime}\succ\mathbf{x}~\textrm{iff}~{f\!\centerdot}^{l}(\mathbf{x}^{\prime})\geq{f\!\centerdot}^{l}(\mathbf{x})~~\forall l\in\{1,\ldots,L\}~\text{and}~\exists l\in\{1,\ldots,L\}~\text{such that}~{f\!\centerdot}^{l}(\mathbf{x}^{\prime})>{f\!\centerdot}^{l}(\mathbf{x}).$ (6)
@@
We define a Pareto set membership (or non-dominance) label $z_{n}:=\mathds{1}[{\mathbf{x}_{n}\in\mathcal{S}_{\textrm{Pareto}}^{t}}]$, where $\mathds{1}[{\cdot}]:\{\text{True},\text{False}\}\to\{1,0\}$, which we will use for training the CPE in active generation. This definition can be extended to the whole domain $\mathcal{X}$ as a labeling function $z(\mathbf{x}):=\mathds{1}[{\mathbf{x}\in\operatorname{Pareto}(\mathcal{S}_{\textrm{Pareto}}^{t}\cup\{\mathbf{x}\})}]$, where $\operatorname{Pareto}(\mathcal{S})$ denotes the Pareto subset of an arbitrary set $\mathcal{S}\subset\mathcal{X}$. Note that the definition remains unchanged for observed points already in the dataset, i.e., $z(\mathbf{x}_{n})=z_{n}$. As it turns out, this non-dominance CPE is also estimating the probability of hypervolume improvement (PHVI) for new query points as the following theorem and corollary show. We fully define hypervolume improvement (HVI) in Sec. B.1, as its definition is a little involved. {restatable}[Equivalence of Indicators]theoremmainthm For every $\mathbf{x}\notin\mathcal{S}_{\textrm{Pareto}}^{t}$, the HVI indicator is equivalent to a non-dominance indicator,
@@
$\mathbb{P}(z(\mathbf{x})=1|\mathbf{x})=\mathbb{P}(\mathrm{HVI}({\mathbf{x}})>0|\mathbf{x}):=\mathrm{PHVI}(\mathbf{x}),\quad\forall\mathbf{x}\notin\mathcal{S}_{\textrm{Pareto}}^{t},$ (9)
as the events are equivalent. Thus, a CPE trained on $z$, using a proper loss, is predicting PHVI.∎ See Sec. B.1 for the proof and assumptions under which this is true. We note that, for existing $\mathbf{x}\in\mathcal{S}_{\textrm{Pareto}}^{t}$, we have a discrepancy, as $z(\mathbf{x})=1$, whereas $\mathrm{HVI}({\mathbf{x}})=0$. Yet, due to observation noise, it is beneficial to allow sampling at $\mathbf{x}$ again for noise reduction (e.g. averaging), as the true ${\mathbf{f}\centerdot}(\mathbf{x})$ may be dominated even if a single observation $\mathbf{y}$ at $\mathbf{x}$ suggests otherwise. In addition, if required, the CPE can be forced to zero at observed points by external mechanisms.
@@ 3 Incorporating User Preferences
where $\mathbf{r}\in\mathbb{R}^{L}$ is a reference point [24, 41, 41, 28]. While scalarization blends objectives according to explicit trade-off weights, it is not optimal when learning a conditional generative model of the Pareto set via a CPE on labels $z_{n}=\mathds{1}[{s_{\boldsymbol{\lambda},\mathbf{r}}(\mathbf{x}_{n})>\tau}]$ for some threshold $\tau\in\real$, as in [10], since each new $\boldsymbol{\lambda}$ would require retraining. Furthermore, thresholding weighted objectives can blur the non-dominance boundary compared to directly labeling it.
@@ 3.1 Preference direction vectors and alignment indicators
Preference directions generalize (convex) scalarization weights to a (non-convex) generative setting: any $\boldsymbol{\lambda}$ can be mapped to a unit-norm vector $\mathbf{u}=\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|_{2}$, and conversely each $\mathbf{u}$ induces a unique normalized weight. By conditioning on $(\mathbf{u},a)$ rather than $\boldsymbol{\lambda}$ alone, our generative Pareto-set model becomes both more flexible (no retraining for new trade-offs) and more faithful to non-dominance structure. We visualize these preference direction vectors in Figure 1.
@@ 4 Amortized Active Generation of Pareto Sets
We now have all the components to describe our amortized active generation framework that learns to generate (approximate) solutions in the Pareto set, conditioned on user preferences. We call our method active generation of Pareto sets (A-GPS), and it begins by generalizing the active generation objective in [35]. That is, for each round, $t$, we minimize the reverse KL divergence between the generative model $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$ and an underlying (unobserved) true model $p({\mathbf{x}}|{\mathbf{u},z,a})$,
@@ 4.1 Estimating A-GPS’s component distributions
We find this occasionally aids exploration. Examples of appropriate parametric forms are von Mises-Fisher distributions, power spherical distributions [11] or normalizing flows [32]. We find Normal distributions, or mixtures, normalized to the unit sphere are more numerically stable than some of the specialized spherical distributions, see Sec. C.1 for more detail. We also find that fitting an unconditional $q_{\gamma}({\mathbf{u}})\approx p({\mathbf{u}})$ for just the initial round can aid exploration of the Pareto front.
@@
As per the original VSD, we define a CPE to directly discriminate over the solution set, $\mathcal{S}$. Sometimes we find setting $\mathcal{S}=\mathcal{S}_{\textrm{Pareto}}^{t}$ leads to overly exploitative behavior. Instead, we anneal the set using the Pareto ranking method described in [12], and define the labels based on a thresholded rank, $k$: $z_{n}=\mathds{1}[{\mathbf{x}_{n}\in\{\bigcup_{k}\mathcal{S}_{\textrm{Pareto}}^{t}{{}_{,k}}:\forall k\leq\tau_{t}\}}]$. Here $k=1$ indicates the Pareto set, $k=2$ the next non-dominated set once $\mathcal{S}_{\textrm{Pareto}}^{t}{{}_{,1}}$ is removed, etc. We use $\tau_{t}=f_{\tau}\!\left({\{\mathbf{y}:\mathbf{y}\in\mathcal{D}_{N}\},\delta_{t}}\right)$ as presented in [35], ensuring $\tau_{T}$ labels just $\mathcal{S}_{\textrm{Pareto}}^{t}$. The annealed method (Equation 20) applied to quantiles of k-k is particularly effective. So, with this label, we use the log-loss to train a CPE,
@@ 4.2 Learning A-GPS’s variational distribution
We find that using ‘on-policy’ gradient estimation methods such as REINFORCE [37, 27] are very slow when we have complex variational distribution forms, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$, e.g. causal transformers. This is because we have to set a low learning rate to avoid the variance of this estimator inducing exploding gradients for long sequences. Also, new samples have to be drawn from the variational distribution every iteration of stochastic gradient descent (SGD), which can be computationally expensive. So instead we use an ‘off-policy’ gradient estimator with importance weights to emulate the on-policy estimator,
@@
Here $w(\mathbf{x},\mathbf{u})=q_{\phi}({\mathbf{x}}|{\mathbf{u}})/q_{\phi^{\prime}}({\mathbf{x}}|{\mathbf{u}})$ are the importance weights [31, 6]. Now we use $S$ samples from $\mathbf{x}^{(s)}\sim q_{\phi^{\prime}}({\mathbf{x}}|{\mathbf{u}^{(s)}})$, to approximate the expectation in Equation 20. If we choose $\phi^{\prime}=\phi$ we recover on-policy gradients, however we typically only update $\phi^{\prime}$ every 100 iterations of optimising A-ELBO, or if the effective sample size drops below a predetermined threshold ($0.5S$). Whenever we update $\phi^{\prime}$ we also resample $S$ samples. We use these estimated gradients with an appropriate SGD algorithm, such as Adam [22], to optimize for $\phi^{*}_{t}$.
@@ 5 Related Work
Traditional MOBO methods, such as hypervolume-based acquisition (EHVI, noisy expected hypervolume improvement (nEHVI), and their variants), entropy search [38, 7, 8, 19] and scalarization methods [24, 41, 28, 10], operate online by sequentially querying the black-box using acquisition rules that balance exploration and exploitation. In contrast, offline MOG approaches like ParetoFlow and guided diffusion frameworks [40, 39, 2] train generative models from a fixed dataset of evaluated designs, without further oracle queries. While these offline methods can leverage rich generative priors, they have not been designed to adapt to new information.
@@
Recent advances in “active generation” recast black-box optimization as fitting conditional generative models to high-value regions, guided by predictors and/or acquisition functions. Methods like VSD, GFlowNets, and diffusion-based solvers [35, 21, 14, 18] show that generative search can match or exceed traditional direct acquisition function optimization, particularly in large search spaces. However, existing generative frameworks often need re-training to integrate subjective preferences. Similarly, all direct acquisition optimization methods require additional optimization runs to incorporate new preferences. An exception is Pareto set learning [25], which learns a neural-net, that maps from scalarization weights to designs.
@@
Or inference-time vs. re-training/fine-tuning based search. Guided generation methods, such as those based on guided diffusion and flow matching [18, 39, 40, 2] use a pre-trained generative model, from which samples are then guided at inference time such that they are generated from a conditional generative model, leaving the original generative model unchanged. It has been noted in [23] that guided methods, though computationally efficient, may be prone to co-variate shift preventing them being guided too far from the support of the pretrained model. Conversely, learning-based methods such as [35, 36] explicitly re-train or fine-tune the generative model to condition it, thereby circumventing these co-variate shift issues at the cost of more computation, but allowing for less constrained exploration in online scenarios.
@@
LaMBO [34] Masked LM nEHVI
LaMBO-2 [18] Diffusion nEHVI
Pareto Set Learning (PSL) [25] Deterministic MLP Scalarization
GFlowNets [21] GFlowNets Scalarization
ParetoFlow [40] Diffusion Scalarization
PROUD [39] Diffusion Multiple grad. desc.
Preference Guided Diffusion [2] Diffusion Preference CPE
CbAS [5] Modular Dominance CPE
VSD [35] Modular Dominance CPE

### v3

@@ 1 Introduction
Traditional Bayesian optimization (BO) methods for black-box multi-objective optimization (MOO) problems, often referred to as multi-objective Bayesian optimization (MOBO), rely on acquisition functions such as expected hypervolume improvement (EHVI) [yang2019efficient, daulton2020differentiable, ament2023unexpected] or their quasi-MonteCarlo extensions [daulton2021parallel, ament2023unexpected], which involve complex numerical integration and/or scale poorly with the number of objectives. Alternatively they may rely on random scalarizations [knowles2006parego, paria2020flexible, de2022mbore] offering simplicity and scalability but rely on sufficient sampling density to capture complex Pareto front geometries.
In this work we propose a fundamentally different approach, in the spirit of multi-objective generation (MOG) [yuan2024paretoflow, yao2024proud], that directly estimates a generative model of the Pareto set in an online sequential, black-box optimization setting. To this end, we build upon the recently proposed variational search distributions (VSD) framework for single-objective optimization [steinberg2025variational]. VSD formulates optimization as active learning of a generative model of high-performing designs (active generation). Thus, VSD alternates between fitting a class probability estimator (CPE) to discriminate favorable designs and updating a conditional generative model to propose new candidates directly. Hence, instead of using a fitness-threshold probability of improvement (PI)-CPE, we propose a Pareto-set CPE, which enables the generative model to focus exclusively on non-dominated designs. This approach bypasses explicit hypervolume computations and scalarizations and leads to a scalable online algorithm for Pareto set approximation that is different from other recent MOG methods [yuan2024paretoflow, yao2024proud]. Furthermore, we show that such a non-dominance CPE is implicitly estimating probability of hypervolume improvement (PHVI). Additionally, practical decision-making often requires incorporating user or stakeholder preferences over the trade-offs among objectives. We introduce a novel mechanism for embedding subjective preference conditioning into the generative model, allowing practitioners to sample candidates that align with specified trade-off directions. By using amortized variational inference (VI) with preference direction vectors, our method supports a-posteriori preference specification without retraining, offering flexibility in downstream design exploration.
@@ 2.1 Active generation
Active generation as implemented by [steinberg2025variational] reframes online black-box optimization as sequential learning of a conditional generative model, guided by a CPE. At each round, $t\in\{1,\dots,T\}$ we: (1) fit a CPE (using some proper loss, $\mathcal{L}_{\text{CPE}}$),
@@
$\theta^{*}_{t}\leftarrow\operatorname*{arg\,\!min}_{\theta}\mathcal{L_{\textrm{CPE}}}\!\left({\theta,\mathcal{D}^{z}_{N}}\right)\qquad\phi^{*}_{t}\leftarrow\operatorname*{arg\,\!max}_{\phi}\mathcal{L_{\textrm{ELBO}}}\!\left({\phi,\theta^{*}_{t}}\right),$ (3)
then samples from $q_{\phi^{*}_{t}}({\mathbf{x}})$ are used to propose new candidates for evaluation. New labels are acquired for these candidates, the dataset is augmented, and the process is repeated until convergence. This solution to active generation is referred to as VSD [steinberg2025variational]. Under certain assumptions on the form of the models, this procedure has proven convergence rates to the ideal $p({\mathbf{x}}|{z=1})$.
@@ 2.2 Optimizing over multiple objectives
$\mathbf{x}^{\prime}\succ\mathbf{x}~\text{iff}~{f\!\centerdot}^{l}(\mathbf{x}^{\prime})\geq{f\!\centerdot}^{l}(\mathbf{x})~~\forall l\in\{1,\ldots,L\}~\text{and}~\exists l\in\{1,\ldots,L\}~\text{such that}~{f\!\centerdot}^{l}(\mathbf{x}^{\prime})>{f\!\centerdot}^{l}(\mathbf{x}).$ (6)
@@
We define a Pareto set membership (or non-dominance) label $z_{n}:=\mathds{1}[{\mathbf{x}_{n}\in\mathcal{S}_{\textrm{Pareto}}^{t}}]$, where $\mathds{1}[{\cdot}]:\{\text{True},\text{False}\}\to\{1,0\}$, which we will use for training the CPE in active generation. This definition can be extended to the whole domain $\mathcal{X}$ as a labeling function $z(\mathbf{x}):=\mathds{1}[{\mathbf{x}\in\operatorname{Pareto}(\mathcal{S}_{\textrm{Pareto}}^{t}\cup\{\mathbf{x}\})}]$, where $\operatorname{Pareto}(\mathcal{S})$ denotes the Pareto subset of an arbitrary set $\mathcal{S}\subset\mathcal{X}$. Note that the definition remains unchanged for observed points already in the dataset, i.e., $z(\mathbf{x}_{n})=z_{n}$. As it turns out, this non-dominance CPE is also estimating the probability of hypervolume improvement (PHVI) for new query points as the following theorem and corollary (assuming $\boldsymbol{\epsilon}=\mathbf{0})$. We fully define hypervolume improvement (HVI) in Sec. B.1, as its definition is a little involved. {restatable}[Equivalence of Indicators]theoremmainthm For every $\mathbf{x}\notin\mathcal{S}_{\textrm{Pareto}}^{t}$, the HVI indicator is equivalent to a non-dominance indicator,
@@
$\mathbb{P}(z(\mathbf{x})=1|\mathbf{x})=\mathbb{P}(\mathrm{HVI}({\mathbf{x}})>0|\mathbf{x})\eqqcolon\mathrm{PHVI}(\mathbf{x}),\quad\forall\mathbf{x}\notin\mathcal{S}_{\textrm{Pareto}}^{t},$ (9)
as the events are equivalent. Thus, a CPE trained on $z$, using a proper loss, is predicting PHVI.∎ See Sec. B.1 for the proof and assumptions under which this is true. We note that, for existing $\mathbf{x}\in\mathcal{S}_{\textrm{Pareto}}^{t}$, we have a discrepancy, as $z(\mathbf{x})=1$, whereas $\mathrm{HVI}({\mathbf{x}})=0$. However, when the objectives are continuous, constructing the indicators from an existing dataset for training a CPE is almost surely equivalent to constructing them in a held-out sense as there will be no ties. Furthermore, if there is observation noise it is beneficial to allow sampling at $\mathbf{x}$ again for noise reduction (e.g. averaging), as the true ${\mathbf{f}\centerdot}(\mathbf{x})$ may be dominated even if a single observation $\mathbf{y}$ at $\mathbf{x}$ suggests otherwise.
@@ 3 Incorporating User Preferences
where $\mathbf{r}\in\mathbb{R}^{L}$ is a reference point [knowles2006parego, zhang2007moea, zhang2007moea, paria2020flexible]. While scalarization blends objectives according to explicit trade-off weights, it is not optimal when learning a conditional generative model of the Pareto set via a CPE on labels $z_{n}=\mathds{1}[{s_{\boldsymbol{\lambda},\mathbf{r}}(\mathbf{x}_{n})>\tau}]$ for some threshold $\tau\in\real$, as in [de2022mbore], since each new $\boldsymbol{\lambda}$ would require retraining. Furthermore, thresholding weighted objectives can blur the non-dominance boundary compared to directly labeling it.
@@ 3.1 Preference direction vectors and alignment indicators
Preference directions generalize (convex) scalarization weights to a (non-convex) generative setting: any $\boldsymbol{\lambda}$ can be mapped to a unit-norm vector $\mathbf{u}=\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|_{2}$, and conversely each $\mathbf{u}$ induces a unique normalized weight. In fact [cheng2016reference, chugh2018surrogate] use “reference vectors”, which are similar to our $\mathbf{u}$, for scalarization of objectives and to promote diversity for many-objective evolutionary algorithms (MOEAs). By conditioning on $(\mathbf{u},a)$ rather than scalarizing by $\boldsymbol{\lambda}$, our generative Pareto-set model becomes both more flexible (no retraining for new trade-offs) and more faithful to non-dominance structure. We visualize these preference direction vectors in Figure 1.
@@ 4 Amortized Active Generation of Pareto Sets
We now have all the components to describe our amortized active generation framework that learns to generate (approximate) solutions in the Pareto set, conditioned on user preferences. We call our method active generation of Pareto sets (A-GPS), and it begins by generalizing the active generation objective in [steinberg2025variational]. That is, for each round, $t$, we minimize the reverse KL divergence between the generative model $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$ and an underlying (unobserved) true model $p({\mathbf{x}}|{\mathbf{u},z,a})$,
@@ 4.1 Estimating A-GPS’s component distributions
We find this occasionally aids exploration. Examples of appropriate parametric forms are von Mises-Fisher distributions, power spherical distributions [de2020power] or normalizing flows [rezende2020normalizing]. We find Normal distributions, or mixtures, normalized to the unit sphere are more numerically stable than some of the specialized spherical distributions, see Sec. C.1 for more detail. We also find that fitting an unconditional $q_{\gamma}({\mathbf{u}})\approx p({\mathbf{u}})$ for just the initial round can aid exploration of the Pareto front.
@@
As per the original VSD, we define a CPE to directly discriminate over the solution set, $\mathcal{S}$. Sometimes we find setting $\mathcal{S}=\mathcal{S}_{\textrm{Pareto}}^{t}$ leads to overly exploitative behavior. Instead, we anneal the set using the Pareto ranking method described in [deb2002fast], and define the labels based on a thresholded rank, $k$: $z_{n}=\mathds{1}[{\mathbf{x}_{n}\in\{\bigcup_{k}\mathcal{S}_{\textrm{Pareto}}^{t}{{}_{,k}}:\forall k\leq\tau_{t}\}}]$. Here $k=1$ indicates the Pareto set, $k=2$ the next non-dominated set once $\mathcal{S}_{\textrm{Pareto}}^{t}{{}_{,1}}$ is removed, etc. We use $\tau_{t}=f_{\tau}\!\left({\{\mathbf{y}:\mathbf{y}\in\mathcal{D}_{N}\},\delta_{t}}\right)$ as presented in [steinberg2025variational], ensuring $\tau_{T}$ labels just $\mathcal{S}_{\textrm{Pareto}}^{t}$. The annealed method (Equation 20) applied to quantiles of k-k is particularly effective. So, with this label, we use the log-loss to train a CPE,
@@ 4.2 Learning A-GPS’s variational distribution
We find that using ‘on-policy’ gradient estimation methods such as REINFORCE [williams1992simple, mohamed2020monte] are very slow when we have complex variational distribution forms, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$, e.g. causal transformers. This is because we have to set a low learning rate to avoid the variance of this estimator inducing exploding gradients for long sequences. Also, new samples have to be drawn from the variational distribution every iteration of stochastic gradient descent (SGD), which can be computationally expensive. So instead we use an ‘off-policy’ gradient estimator with importance weights to emulate the on-policy estimator,
@@
Here $w(\mathbf{x},\mathbf{u})=q_{\phi}({\mathbf{x}}|{\mathbf{u}})/q_{\phi^{\prime}}({\mathbf{x}}|{\mathbf{u}})$ are the importance weights [precup2000eligibility, burda2016importance]. Now we use $S$ samples from $\mathbf{x}^{(s)}\sim q_{\phi^{\prime}}({\mathbf{x}}|{\mathbf{u}^{(s)}})$, to approximate the expectation in Equation 20. If we choose $\phi^{\prime}=\phi$ we recover on-policy gradients, however we typically only update $\phi^{\prime}$ every 100 iterations of optimising A-ELBO, or if the effective sample size drops below a predetermined threshold ($0.33S$). Whenever we update $\phi^{\prime}$ we also resample $S$ samples. We use these estimated gradients with an appropriate SGD algorithm, such as Adam [kingma2014adam], to optimize for $\phi^{*}_{t}$.
@@ 5 Related Work
Traditional MOBO methods, such as hypervolume-based acquisition (EHVI, noisy expected hypervolume improvement (nEHVI), and their variants), entropy search [yang2019efficient, daulton2020differentiable, daulton2021parallel, hernandez2016predictive] and scalarization methods [knowles2006parego, zhang2007moea, paria2020flexible, de2022mbore], operate online by sequentially querying the black-box using acquisition rules that balance exploration and exploitation. In contrast, offline MOG approaches like ParetoFlow and guided diffusion frameworks [yuan2024paretoflow, yao2024proud, annadani2025preference] train generative models from a fixed dataset of evaluated designs, without further oracle queries. While these offline methods can leverage rich generative priors, they have not been designed to adapt to new information.
@@
Recent advances in “active generation” recast black-box optimization as fitting conditional generative models to high-value regions, guided by predictors and/or acquisition functions. Methods like VSD, GFlowNets, and diffusion-based solvers [steinberg2025variational, jain2023multi, dhariwal2021diffusion, gruver2023protein] show that generative search can match or exceed traditional direct acquisition function optimization, particularly in large search spaces. However, existing generative frameworks often need re-training to integrate subjective preferences. Similarly, all direct acquisition optimization methods require additional optimization runs to incorporate new preferences. An exception is Pareto set learning [lin2022pareto], which learns a neural-net, that maps from scalarization weights to designs.
@@
Or inference-time vs. re-training/fine-tuning based search. Guided generation methods, such as those based on guided diffusion and flow matching [gruver2023protein, yao2024proud, yuan2024paretoflow, annadani2025preference] use a pre-trained generative model, from which samples are then guided at inference time such that they are generated from a conditional generative model, leaving the original generative model unchanged. It has been noted in [klarner2024context] that guided methods, though computationally efficient, may be prone to co-variate shift preventing them being guided too far from the support of the pretrained model. Conversely, learning-based methods such as [steinberg2025variational, wang2025finetuning] explicitly re-train or fine-tune the generative model to condition it, thereby circumventing these co-variate shift issues at the cost of more computation, but allowing for less constrained exploration in online scenarios.
@@
LaMBO [stanton2022accelerating] Masked LM nEHVI
LaMBO-2 [gruver2023protein] Diffusion nEHVI
Pareto Set Learning (PSL) [lin2022pareto] Deterministic MLP Scalarization
GFlowNets [jain2023multi] GFlowNets Scalarization
ParetoFlow [yuan2024paretoflow] Diffusion Scalarization
PROUD [yao2024proud] Diffusion Multiple grad. desc.
Preference Guided Diffusion [annadani2025preference] Diffusion Preference CPE
CbAS [brookes2019conditioning] Modular Dominance CPE
VSD [steinberg2025variational] Modular Dominance CPE